OpenStudy (anonymous):

The area of a rectangular piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

OpenStudy (anonymous):

PLEASE HELP ILL MEDAL

OpenStudy (tkhunny):

W = Width of Rectangle L = Length of Rectangle L*W = 280 Okay, now what?

OpenStudy (anonymous):

I already have part A. just not sure about part B.

OpenStudy (anonymous):

Part A: 280 = (l-5)(w+1)

OpenStudy (tkhunny):

Show your work, please. 1) It would be VERY helpful for you to say what you need. 2) It would be WONDERFULLY helpful if you would show your work on the first post.

OpenStudy (anonymous):

i just used the formula like you posted.

OpenStudy (anonymous):

i used mathway.com to help me with Part B, and i got 285-l/l-5

OpenStudy (anonymous):

\[\frac{ 285-l }{ l-5 }\]

OpenStudy (tkhunny):

The hint about the Square is only to give us a hint about he Length and Width. Those adjustments do NOT tell us much about the rectangle. We have L*W = 280. We have L-5 = W+1 from the additional information.

OpenStudy (tkhunny):

There are two way sot go to create an equation for finding the Width. L = 280/W ==> (280/W) - 5 = W+1 Or L = W + 6 ==> (W+6)W = 280 They should lead to the same conclusion about W.

OpenStudy (tkhunny):

* ways to go

OpenStudy (anonymous):

I got it figured out, Thank you so much!

OpenStudy (triciaal):

so what answer did you get?

OpenStudy (anonymous):

Part A: 280 = (L-5)(x+1) Part B: x = 285−L/L−5

OpenStudy (triciaal):

x? same as w right what numbers for the measurements

OpenStudy (anonymous):

yes x=w

OpenStudy (tkhunny):

No. 280 = (L-5)(X+1) is not meaningful. We are NOT given the area of the imaginary square. 280 is the area of the rectangle. We are NOT told that the square and the rectangle have the same area.

OpenStudy (tkhunny):

Given (W+6)W = 280, we have W^2 + 6W - 280 = 0 and this has solutions, W = 14 or W = -20. As W = -20 is not reasonable, we conclude that W = 14. Also, not that x = 285−L/L−5 is NOT what you intend. Parentheses are NOT optional in this form, If you meant x = (285−L)/(L−5), that would be closer, but this does not satisfy the requirements of the problem statement.

OpenStudy (triciaal):

and L = 20

OpenStudy (tkhunny):

L = 280 / W = 280 / 14 = 20 -- Very good. Note: (L-5) = 15 (W+1) = 15 They are equal, thus a square, but 15*15 = 225 and that is not 280.

OpenStudy (triciaal):

I agree !!! if it were a square it would be 15 *15 but because it is a rectangle it is 14 *20

OpenStudy (triciaal):

that's why they gave us the relationship between the 2 sets of measurements